# PEAR: Equal Area Weather Forecasting on the Sphere

### v1

@@ 6 Limitations
Our limited compute budget restricts us to ERA5-lite, making comparisons to models trained on the full ERA5 dataset harder. Ideally we would train on ground truth data created on the HEALPix grid, but at this time the reanalysis for ERA5 is done on the Driscoll-Healy grid.
@@ 7 Conclusion
We have shown the importance of using the HEALPix grid that minimizes unphysical biases for global medium term weather prediction. The hierarchical equal area pixelation enables efficient implementation of our transformer architecture PEAR, that outperforms its counterpart on the traditional Driscoll-Healy grid at no computational overhead. With forecast horizons of up to 10 days, we showed that PEAR outperforms the Driscoll-Healy baseline with more than twice the number of trainable parameters. The superior performance of PEAR should also lead to more accurate extreme weather forecasting [olivetti2024], a direction that would be interesting for future work.
@@ Appendix A Appendix
We include anomaly correlation coefficient (ACC) and root mean squared error (RMSE) for all upper variables and pressure levels in Fig. 8 and Fig. 9 respectively. PEAR outperforms Pangu for all variables and pressure levels for ACC in Fig. 8, and often outperforms the almost eight times larger PanguLarge.
We also include spatially resolved RMSE over the validation year for PEAR in Fig. 6 (surface) and 10 (upper), as well as the corresponding spatial RMSE for Pangu in Fig. 7 (surface) and Fig.11 (upper).
Figure 6: Spatial RMSE for PEAR predictions of surface variables with one day lead time averaged over the validation year. The HEALPix predictions are projected to a cartesian longitude (horizontal) and latitude (vertical) grid for visualization.
Figure 7: Spatial RMSE for Pangu predictions of surface variables with one day lead time averaged over the validation year.
Figure 8: Anomaly correlation coefficient (ACC) for the upper variables separated by pressure level over up to 10 days lead time.
Figure 9: Root mean squared error (RMSE) for the upper variables separated by pressure level over up to 10 days lead time.
Figure 10: Spatial RMSE for PEAR predictions of upper variables with one day lead time averaged over the validation year. Pressure level (rows) and variables (columns) with joint color mapping per variable. The HEALPix predictions are projected to a cartesian longitude (horizontal) and latitude (vertical) grid for visualization.
Figure 11: Spatial RMSE for Pangu predictions of upper variables with one day lead time averaged over the validation year. Pressure level (rows) and variables (columns) with joint color mapping per variable.

### v2

@@ 6 Limitations
Our limited compute budget restricts us to ERA5-lite, making comparisons to models trained on the full ERA5 dataset harder. Ideally we would train on ground truth data created on the HEALPix grid, but at this time the reanalysis for ERA5 is done on an equiangular grid.
@@ 7 Conclusion
We have shown the importance of using the HEALPix grid that minimizes unphysical biases for global medium term weather prediction. The hierarchical equal area pixelation enables efficient implementation of our transformer architecture PEAR, that outperforms its counterpart on the traditional equiangular grid at no computational overhead. With forecast horizons of up to 10 days, we showed that PEAR outperforms the equiangular baseline with more than twice the number of trainable parameters. The superior performance of PEAR should also lead to more accurate extreme weather forecasting [olivetti2024], a direction that would be interesting for future work.
@@ Appendix A Appendix
Table A1: PEAR architecture overview. The windowed attention blocks contain multiple copies, indicated by the depth column, of the multi-head attention layer and layer normalization. The tensor shape column shows the shape of the output tensor from the corresponding layer. Both the input and final patch recovery layer use two separate tensors for the surface and upper variables.
Nr Layer/Block Depth Tensor shape Attention heads
1 Patch embed 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$
2 Windowed attention 2 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$ 6
3 Downsample 1 $(\nicefrac{{3}}{{16}}n_{\mathrm{side}}^{2},8,96)$
4 Windowed attention 12 $(\nicefrac{{3}}{{16}}n_{\mathrm{side}}^{2},8,96)$ 12
5 Upsample 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$
6 Windowed attention 2 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$ 6
Concatenate 6 & 2 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,96)$
Figure A1: Patches (green) and windows (red) at the first attention layer. Each green patch contains four pixels of the input resolution $n_{\mathrm{side}}=64$ on the HEALPix grid, and each window contains 64 patches at this level. The window also include one vertical level above, not shown in the figure.
An overview of the architecture of PEAR can be found in Table. A1. The patches and windows are visualized in Figure. A1.
We use RMSE and Anomaly Correlation Coefficient (ACC) calculated on the HEALPix grid according to
where $\Delta y$ is the difference between the predictions and the climatology average.
We include anomaly correlation coefficient (ACC) and root mean squared error (RMSE) for all upper variables and pressure levels in Fig. A4 and Fig. A5 respectively. PEAR outperforms Pangu for all variables and pressure levels for ACC in Fig. A4, and often outperforms the almost eight times larger PanguLarge.
We also include spatially resolved RMSE over the validation year for PEAR in Fig. A2 (surface) and A6 (upper), as well as the corresponding spatial RMSE for Pangu in Fig. A3 (surface) and Fig.A7 (upper).
Figure A2: Spatial RMSE for PEAR predictions of surface variables with one day lead time averaged over the validation year. The HEALPix predictions are projected to a cartesian longitude (horizontal) and latitude (vertical) grid for visualization.
Figure A3: Spatial RMSE for Pangu predictions of surface variables with one day lead time averaged over the validation year.
Figure A4: Anomaly correlation coefficient (ACC) for the upper variables separated by pressure level over up to 10 days lead time.
Figure A5: Root mean squared error (RMSE) for the upper variables separated by pressure level over up to 10 days lead time.
Figure A6: Spatial RMSE for PEAR predictions of upper variables with one day lead time averaged over the validation year. Pressure level (rows) and variables (columns) with joint color mapping per variable. The HEALPix predictions are projected to a cartesian longitude (horizontal) and latitude (vertical) grid for visualization.
Figure A7: Spatial RMSE for Pangu predictions of upper variables with one day lead time averaged over the validation year. Pressure level (rows) and variables (columns) with joint color mapping per variable.